### v1

# VIST3A: Text-to-3D by Stitching a Multi-view Reconstruction Network to a Video Generator

Abstract
The rapid progress of large, pretrained models for both visual content generation and 3D reconstruction opens up new possibilities for text-to-3D generation. Intuitively, one could obtain a formidable 3D scene generator if one were able to combine the power of a modern latent text-to-video model as “generator” with the geometric abilities of a recent (feedforward) 3D reconstruction system as “decoder”. We introduce VIST3A, a general framework that does just that, addressing two main challenges. First, the two components must be joined in a way that preserves the rich knowledge encoded in their weights. We revisit model stitching, i.e., we identify the layer in the 3D decoder that best matches the latent representation produced by the text-to-video generator and stitch the two parts together. That operation requires only a small dataset and no labels. Second, the text-to-video generator must be aligned with the stitched 3D decoder, to ensure that the generated latents are decodable into consistent, perceptually convincing 3D scene geometry. To that end, we adapt direct reward finetuning, a popular technique for human preference alignment. We evaluate the proposed VIST3A approach with different video generators and 3D reconstruction models. All tested pairings markedly improve over prior text-to-3D models that output Gaussian splats. Moreover, by choosing a suitable 3D base model, VIST3A also enables high-quality text-to-pointmap generation.
Figure 1: Text-to-3D generation with VIST3A. Video models excel at generating latent visual content from text prompts, whereas 3D foundation models shine when it comes to decoding such a latent representation into consistent scene geometry. By stitching a video generator and a 3D reconstruction network together and aligning their latents, we obtain an end-to-end model that produces high-quality Gaussian splats (a) or point maps (b) from text prompts.
footnotetext: Project page: https://gohyojun15.github.io/VIST3A/
1 Introduction
With image and video generators now a commodity, text-to-3D models that produce 3D scenes from text prompts have become a new research frontier, with applications in AR/VR, gaming, robotics, and simulation. Early methods for 3D generation adopt Score Distillation Sampling (SDS) (Poole et al., 2023; Tang et al., 2024b; Wang et al., 2023b; Chen et al., 2024b) to optimize a 3D representation, e.g. a NeRF (Mildenhall et al., 2021; Müller et al., 2022) or 3D Gaussian Splats (3DGS, Kerbl et al., 2023) under a pretrained 2D diffusion prior (Rombach et al., 2022). A drawback these methods have in common is the need for slow per-scene optimization. Another line of work uses multi-stage pipelines that first synthesize images and then lift them to 3D with a separate model (Tang et al., 2024a; Xu et al., 2024b; Zhang et al., 2024b) or with per-scene optimization (Gao et al., 2024; Wu et al., 2024a; Yu et al., 2024b); employ progressive warping and refinement (Shriram et al., 2025; Yu et al., 2025; 2024a); or sequentially chain multiple generative modules (Yang et al., 2025b; Engstler et al., 2025). The multi-stage design not only increases model complexity and engineering effort, but also makes such models prone to error accumulation (Lin et al., 2025; Meng et al., 2025).
Figure 2: Comparison with existing, LDM-based 3D generators. Instead of training a custom decoder from multi-view 2D latents to 3D outputs, we stitch and align an existing, pretrained 3D reconstruction model.
A recent trend is to directly generate the 3D representation with end-to-end latent diffusion models  (LDMs, Schwarz et al., 2025; Lan et al., 2024; Li et al., 2025b; a; Bahmani et al., 2025). A prominent line of work starts from pretrained 2D image (Esser et al., 2024; Rombach et al., 2022) or video models (Genmo Team, 2024; Yang et al., 2024b) and finetunes them to output multi-view 2D latents, reusing the pretrained priors (Szymanowicz et al., 2025; Liang et al., 2025; Schwarz et al., 2025; Lin et al., 2025; Yang et al., 2025c; Go et al., 2025a; b). Subsequently, a VAE-style decoder is trained to decode those latents into the desired 3D representation, see Fig. 2. The LDM-like design unifies 2D generation and multi-view reconstruction within the latent space and enables efficient 3D scene generation with a compact, well-amortized decoder.
Still, two key limitations remain. First, we argue that the Achilles heel of existing 2D-to-3D diffusion models is the decoder. By simply repurposing the 2D VAE to produce 3D outputs, the network must learn 3D reconstruction more or less from scratch, which requires extensive training and large datasets that are hard to obtain (Yang et al., 2025c; Szymanowicz et al., 2025; Go et al., 2025b). This practice becomes increasingly problematic as new, better 3D foundation models emerge  (Wang et al., 2025d; a; 2024b; Zhang et al., 2025) and the ad-hoc trained decoders of text-to-3D models fall further behind the state of the art in 3D vision.
Second, the prevalent training scheme tends to suffer from weak alignment between the generative model and the VAE decoder. Typically, the former is finetuned on multi-view datasets with a generative objective like a diffusion loss (Song et al., 2020; Sohl-Dickstein et al., 2015; Ho et al., 2020) or flow matching (Liu et al., 2023; Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023), which only indirectly promotes 3D-consistent latents. Moreover, the separate training may cause the latents, even if 3D-consistent, to be out of domain from the perspective of the decoder. To mitigate that misalignment, it has been proposed to add rendering losses that promote decodable latents (Lin et al., 2025). However, the resulting objective is based on single-step sampling and does not sufficiently take into account the denoising trajectory, leading to weak alignment at inference.
We introduce VIST3A: VIdeo VAE STitching and 3D Alignment. The proposed method consists of two complementary components that address the above-mentioned limitations, see Fig. 2. First, we resort to the concept of model stitching (Pan et al., 2023; Lenc & Vedaldi, 2015; Bansal et al., 2021; Csiszárik et al., 2021; Yang et al., 2022) to leverage powerful, pretrained feedforward 3D models for decoding, rather than start from scratch. The idea is to attach the relevant part of a 3D reconstruction network as a “decoder” to the latent space of a video VAE. For this to work, there need to be one or more layers in the 3D model whose activations are similar (up to a linear transformation) to those in the VAE’s latent space, despite their independent pretraining. Perhaps surprisingly, this turns out to be the case. For the 3D model, we identify the layer with the most linear relation to the LDM latents, slice the network before that layer, and retain the downstream portion as 3D decoder. After fitting a single, linear stitching layer (in closed form), the VAE latent space already matches the expected input of the 3D decoder well, such that subsequent fine-tuning will be minor and not degrade the respective generative and 3D reasoning capabilities of the two base models.
Second, we further improve alignment between the generative model and the stitched decoder through direct reward finetuning (Clark et al., 2023; Xu et al., 2023; Prabhudesai et al., 2024; Wu et al., 2024c; Shen et al., 2025). In that technique, commonly used to align diffusion models with human preferences, reward signals are defined based on the “goodness” of the VAE output – in our setting, the visual quality and 3D consistency of the decoded 3D representations. Maximizing these rewards encourages the LDM to produce latents that are 3D-consistent and lie within the decoder’s input domain, ensuring high-quality outputs. Importantly, our alignment compares video model outputs and images rendered from the generated 3D scenes, hence it does not require labels.
In our experiments, we show that the proposed stitching scheme is applicable across a range of video generative models and also across several different feedforward 3D models. VIST3A’s direct 3D decoding consistently outperforms prior text-to-3DGS methods, and additionally offers high-quality pointmap generation from text prompts.
2 Related Works
3D generation. Recent works have explored various 3D representations for generative modelling, including point clouds (Mo et al., 2023; Nichol et al., 2022; Vahdat et al., 2022), meshes (Xu et al., 2024a; Woo et al., 2024), voxel grids (Sanghi et al., 2023), NeRFs (Chen et al., 2023; Müller et al., 2022; Mildenhall et al., 2021), and 3DGS (Henderson et al., 2024; Zhang et al., 2024a; Kerbl et al., 2023). Score distillation using 2D diffusion models is time-consuming, as it requires per-scene test time optimization (Wang et al., 2023a; Shi et al., 2023; Wang et al., 2023b), while multi-stage pipelines (Yu et al., 2024b; Liu et al., 2024; Zheng et al., 2025) lack robustness and create significant engineering overhead. For further details on multi-stage pipelines, please refer to Appendix A.
More recently, the field has shifted towards end-to-end latent diffusion models, where the generator operates in the latent space of a VAE, and the latter directly decodes the resulting latents to 3D outputs. Many of these works focus on object-centric asset generation (Wu et al., 2024b; Zhao et al., 2023; Lin et al., 2025) and train the LDM on curated datasets such as Objaverse (Deitke et al., 2023), with single objects or bounded scenes, and controlled camera paths. Consequently, they are unable to handle real-world challenges like strongly varying scene scale, variable lighting, etc.
To tackle such situations, recent methods (Szymanowicz et al., 2025; Liang et al., 2025; Schwarz et al., 2025; Lin et al., 2025; Yang et al., 2025c; Go et al., 2025a; b; Bahmani et al., 2025) repurpose the comprehensive knowledge of the visual world that is implicit in 2D image generators. The general strategy is to finetune a pretrained 2D model on multi-view data, by using generative losses to enforce cross-view consistency. In many cases training is further supported by additional 3D cues like camera poses (Li et al., 2024; Go et al., 2025b), depthmaps (Go et al., 2025a; Yang et al., 2025c), or pointmaps (Szymanowicz et al., 2025). The resulting multi-view latents are decoded to 3D scenes with a dedicated VAE-style decoder, meaning that 3D reasoning capabilities must be rebuilt from scratch, and that they are only weakly aligned with the generator output – limitations which we address with VIST3A.
Learned 3D reconstruction. A notable trend in 3D computer vision is the trend to move away from multi-stage pipelines and iterative optimization towards end-to-end, feedforward 3D modelling. Classical reconstruction pipelines based on SfM (Hartley & Zisserman, 2003; Schönberger & Frahm, 2016) and MVS (Furukawa et al., 2015; Schönberger et al., 2016) require incremental, iterative optimization, whereas recent advances like DUSt3R (Wang et al., 2024b) and MASt3R (Leroy et al., 2024) directly predict 3D point maps in one forward pass. Several follow-up works have further reduced test-time optimization (Tang et al., 2025; Wang et al., 2025b; Yang et al., 2025a). Likewise, 3D Gaussian splatting has evolved from per-scene optimization to feedforward prediction (Charatan et al., 2024; Chen et al., 2024a; Ye et al., 2024). Once more, data scaling has been a critical factor (Wang et al., 2025a; d). Consequently, replicating the 3D capabilities of recent feedforward models as part of VAE training would be difficult and costly. VIST3A offers a solution by reusing, rather than rebuilding, models like AnySplat (Jiang et al., 2025), VGGT (Wang et al., 2025a), or MVDUSt3R (Tang et al., 2025).
Model stitching. Recomposing the heads and tails of two different networks was initially studied as a way to assess the equivariance of neural representations (Lenc & Vedaldi, 2015), and as an experimental tool to compare two different representations (Csiszárik et al., 2021; Bansal et al., 2021). To ensure invariance against trivial affine transformations, the head of some trained network $A$ is normally attached to the tail of another network $B$ via a linear, trainable stitching layer. Besides revealing similarities between networks that common metrics like CKA (Kornblith et al., 2019) would miss, it was also found that different architectures that were trained on the same data can often be stitched into a new, hybrid model with minimal degradation (Bansal et al., 2021). This has opened the door for practical uses of stitching, e.g. DeRy (Yang et al., 2022) for resource-constrained reassembly of pretrained models and SN-Net (Pan et al., 2023) to build networks with varying scales. Going one step further, we demonstrate that strong 3D VAEs [^1] To be consistent with existing literature (Lan et al., 2024; Yang et al., 2025c), we also use the term “3D VAE”, although the mapping from 2D images to 3D scene is, technically, not a variational auto-encoder. can be obtained by stitching a foundational 3D model to the latent space of a video VAE as its decoder, even if they were trained independently on different data.
3 Methodology
VIST3A consists of two key components, see Fig. 3: (1) model stitching to optimally attach (part of) a foundational 3D model as the decoder for the latent, and (2) direct reward finetuning to optimize the alignment of the (latent) generative model with that new decoder.
Figure 3: VIST3A constructs a 3D VAE through model stitching (top), then aligns it with a generative model via direct reward finetuning (bottom). Stitching repurposes a part of a pretrained 3D vision model as decoder to obtain a 3D VAE. Direct reward finetuning simulates full-trajectory denoising, forcing the generative model to produce 3D-consistent, decodable latents.
3.1 Model Stitching for 3D VAE Construction
Our objective is to build a 3D VAE by seamlessly combining the encoder of a video LDM and a feedforward 3D reconstruction model. Note that, for stitching purposes, one can skip the denoising loop, since feeding images into the encoder already gives clean latents. Let $\mathcal{E}$ denote the encoder and $\mathcal{D}$ the decoder of the VAE, and let $F_{1:l}({\bm{x}})=f_{l}\circ\cdots\circ f_{1}({\bm{x}})={\bm{y}}$ be the feedforward 3D network that maps a set of views $\bm{x}$} to a 3D output $\bm{y}$}, with $l$ the total number of layers in that feedforward model. As shown in Fig. 3, we cut the feedforward model at layer $k^{*}$ and stitch the downstream part $F_{k^{\star}+1:l}=f_{l}\circ\cdots\circ f_{k^{\star}+1}$ to the output layer of the encoder $\mathcal{E}$, with the help of a linear stitching layer $\mathbf{S}$. In doing so, we obtain a new 3D VAE $\mathcal{M}_{\text{stitched}}$ that outputs the same representation $\hat{{\bm{y}}}$ as the original 3D model:
$\mathcal{M}_{\text{stitched}}=F_{k^{\star}+1:l}\circ\mathbf{S}\circ\mathcal{E}({\bm{x}})=\hat{{\bm{y}}},\quad\mathcal{D}_{\text{stitched}}=F_{k^{\star}+1:l}\circ\mathbf{S}$ (1)
The front portion $F_{1:k^{\star}}$ of the 3D model is discarded – but if the clean encoder latents, after the affine warping $\mathbf{S}$, are (almost) the same as the activations $f_{k^{\star}}$, then the back portion will still produce the same output, $\hat{{\bm{y}}}\approx{\bm{y}}$. In other words, the stitched VAE $\mathcal{M}_{\text{stitched}}$ is an approximation of the original 3D model $F$. It retains much of the ability to map multi-view images to a 3D reconstruction and only requires a little fine-tuning to restore that ability.
Step 1: Finding the stitching index and initialization. To identify the layer $k^{*}$ in the 3D model whose representation is most compatible with the VAE latent, we first push a set of $N$ samples through the encoder $\mathcal{E}$ to obtain their latents $\mathbf{B}\in\mathbb{R}^{N\times D_{\mathcal{E}}}$. Then, we scan over candidate layers $k\in\{1,...,l-1\}$ of the 3D model and, for each layer in turn, extract the activations $\mathbf{A}_{k}\in\mathbb{R}^{N\times D_{F}}$ and fit the linear stitching layer $\mathbf{S^{*}}_{k}\in\mathbb{R}^{D_{\mathcal{E}}\times D_{F}}$ that best recovers the activations of the 3D model at layer $k$, by solving a least-squares problem:
$\mathbf{S^{*}}_{k}=\arg\min_{\mathbf{S}_{k}}\|\mathbf{B}\mathbf{S}_{k}-\mathbf{A}_{k}\|_{F}^{2}=\big(\mathbf{B}^{\top}\mathbf{B}\big)^{-1}\mathbf{B}^{\top}\mathbf{A}_{k}.\vskip-5.69046pt$ (2)
Finally, we select the stitching layer $k^{\star}$ that leads to the smallest (mean squared) error, $k^{\star}=\arg\min_{k}\|\mathbf{B}\mathbf{S}^{*}_{k}-\mathbf{A}_{k}\|_{F}^{2},$ and assemble the 3D VAE by concatenating $\mathcal{E}$, $\mathbf{S}^{*}_{k^{\star}}$ and $F_{k^{\star}+1:l}$. Empirically, we find that most combinations of foundational VAEs and 3D feedforward models can be stitched in this manner, with minimal performance loss.
Step 2: Stitched decoder finetuning. To further reduce the remaining discrepancies between the newly assembled 3D VAE and the original 3D model, we finetune $\mathbf{S}$ and $F_{k^{\star}+1:l}$ to reproduce the predictions of the original 3D model $\bm{y}$}, using them as pseudo-targets. Practical feedforward models produce multiple outputs (e.g., point maps, depth, poses), so we optimize a weighted sum of $\ell_{1}$ losses for all of them. Note that the fine-tuning step is self-supervised and does not require labels. In our implementation, we restrict the stitching layer to a 3D convolution and employ LoRA (Hu et al., 2022) for updating $F_{k^{\star}+1:l}$, to prevent large deviations from the pretrained weights. For further details, see Appendix B.1.
3.2 Alignment via Direct Reward Finetuning
So far, we have assembled a 3D VAE with a strong, pretrained 3D decoder. However, during text-to-3D inference, the latents are not obtained from the encoder but generated from noise by the denoising loop conditioned on the text prompt. Therefore, we must also align the generative model itself with the 3D decoder, such that it produces decodable latents.
Previous work finetunes the generative network by minimizing generative losses over some multi-view dataset. Unfortunately, that strategy does not ensure 3D-consistent latents. Even if it did, the finetuning bypasses the decoder, hence there is no guarantee that the generated latents fall within the distribution expected by the 3D VAE and can be decoded to meaningful outputs.
To address the disconnect between the denosing loop and the 3D VAE, we adopt direct reward finetuning to align the two. In other words, we extend conventional, generative multi-view finetuning with reward maximization. The conventional generative loss $L_{\text{gen}}$ uses paired data, i.e., multi-view images and corresponding prompts. In contrast, the proposed reward term $r(\cdot,c)$ relies only on the text prompt and requires no ground-truth images. Our total loss is defined as
$L_{\text{total}}=L_{\text{gen}}-r\big(z_{0}(\theta,c,z_{T}),c\big),$ (3)
where $\theta$ are the parameters of the video generative model, $c$ represents the text prompt, $z_{T}$ is the initial noise, and $z_{0}(\theta,c,z_{T})$ is the final latent produced by the denoising loop.
Reward. The proposed reward function consists of three components that ensure high-quality and 3D-consistent generation. (1) Multi-view Image Quality: As we keep the encoder frozen, the generated latents can be decoded by the original video decoder $\mathcal{D}$ to obtain multi-view images. We evaluate these images against the input prompt using CLIP-based (Fang et al., 2024) and HPSv2 human preference scores (Wu et al., 2023) to promote prompt adherence and visual quality, similar to DanceGRPO (Xue et al., 2025). (2) 3D Representation Quality: To encourage high-quality 3D outputs after decoding with $\mathcal{D}_{\text{stitched}}$, we render the generated 3D scenes (pointmaps and/or 3DGS) back into 2D views and apply the same (CLIP + HPSv2) metrics to them as above. (3) 3D Consistency: To enforce 3D consistency, we render the 3D representation from the same viewpoints as the multi-view images reconstructed by the video decoder $\mathcal{D}$, using the camera poses predicted by the feedforward 3D model. We then compute a combination of $\ell_{1}$-loss and LPIPS (Zhang et al., 2018) for each pair of decoded and rendered images belonging to the same viewpoint. The final (negative) reward is a weighted sum of these three losses. For further details, see Appendix B.2.
Alignment algorithm. To optimize the generative model according to the reward function above, we employ direct reward finetuning (Clark et al., 2023; Xu et al., 2023; Prabhudesai et al., 2024; Wu et al., 2024c; Shen et al., 2025). I.e., the model generates samples by unfolding the full denoising path, and the rewards computed from these samples are then backpropagated through the denoising chain. While the algorithm benefits from gradient-based feedback, it can also suffer from exploding gradient norms. To stabilize the optimization, we generalize the idea of DRTune (Wu et al., 2024c): gradients are detached from the inputs to the generative model, but retained during the update step to the next denoising state. In this way, reward propagation remains stable even at early denoising steps. Furthermore, we modify the optimizer for better computational efficiency by (i) randomized sampling, using fewer timesteps than during inference, and (ii) randomizing the subset of denoising steps where gradients are backpropagated, such that the model learns from diverse denoising trajectories. For further details, see Appendix B.2.
In summary, we perform joint, end-to-end alignment of the VAE and the generative model, unlike conventional multi-view fine-tuning that keeps them separate. Reward tuning ensures that, throughout the iterative denoising process, the generative model remains aligned with our 3D VAE and generates latents that suit the stitched decoder.
4 Experimental Results
In what follows, we demonstrate VIST3A’s text-to-3D generation performance. The main findings are that VIST3A clearly outperforms existing feedforward text-to-3DGS approaches and also offers high-quality text-to-pointmap generation. Moreover, we experimentally analyze our two core components, self-supervised model stitching and alignment finetuning.
4.1 Experimental Setups
We provide a high-level overview of the experimental setup. A complete description of evaluation protocols and training details can be found in Appendix C.
Target 3D models. We target last-generation foundational 3D vision models that have been trained on large-scale datasets, have demonstrated generality and reliable performance across diverse domains, and require only images as input. For our experiments, we select three representative state-of-the-art models: (1) MVDUSt3R (Tang et al., 2025) predicts pointmaps and Gaussian splats, (2) VGGT (Wang et al., 2025a) predicts pointmaps, depth maps and camera poses, and (3) AnySplat (Jiang et al., 2025) predicts Gaussian splats and camera poses.
Target video generators. Our primary video model is Wan 2.1 T2V large (Wan et al., 2025), a state-of-the-art text-to-video generator. To demonstrate the generality of VIST3A across different architectures, we additionally use several other latent video models, including CogVideoX (Yang et al., 2024b), SVD (Blattmann et al., 2023), and HunyuanVideo (Kong et al., 2024).
Training data. We finetune stitched VAEs on DL3DV-10K (Ling et al., 2024) and ScanNet (Dai et al., 2017), without 3D labels. To align the video generator in latent space, we utilize DL3DV-10K to compute the generative loss, with prompts from the HPSv2 training set (Wu et al., 2023).
4.2 Main Results: 3D Generation
Stitching Wan to the 3D models listed in Section 4.1 yields two types of generative models: (i) Text-to-3DGS when using AnySplat or MVDUSt3R as decoder; and (ii) Text-to-Pointmap when using VGGT or MVDUSt3R. Both variants are evaluated in the following.
Baselines. Important baselines for text-to-3DGS are SplatFlow (Go et al., 2025a), Director3D (Li et al., 2024), Prometheus3D (Yang et al., 2025c), and VideoRFSplat (Go et al., 2025b). Additionally, we include Matrix3D-omni (Yang et al., 2025d), to our knowledge, the only other model that unifies generation and reconstruction in latent space.
Evaluation protocol. We evaluate text-to-3DGS models on three benchmarks: T3bench (He et al., 2023) for object-centric generation, SceneBench (Yang et al., 2025c) for scene-level synthesis, and DPG-bench (Hu et al., 2024) to assess adherence to long, detailed prompts. On T3bench and SceneBench, we render images and compute Imaging Quality and Aesthetic Quality scores as defined by VBench (Huang et al., 2024) to assess visual fidelity, CLIP score (Hessel et al., 2021) for text-prompt alignment, and Alignment, Coherence, and Style scores according to Wang et al. (2025c) as comprehensive quality metrics. We prefer to avoid traditional no-reference metrics like NIQE (Mittal et al., 2012b) and BRISQUE (Mittal et al., 2012a) that have sometimes been used in the context of 3D generation, but lack a meaningful connection to the conditional generation task (e.g., they can be gambled by always returning the same sharp and colorful, high-scoring image, independent of the prompt). For DPG-bench, we follow the suggested protocol (Hu et al., 2024), but upgrade from the originally proposed language models to the more capable, UnifiedReward LLM (based on Qwen 7B). Text-to-pointmap models are evaluated qualitatively, as no established benchmarks or baselines exist.
Table 1: Quantitative results on T3Bench and SceneBench.
Method T3Bench (Object-centric) SceneBench (Scene-level)
[…]
Matrix3D-omni 43.05 37.66 25.06 2.44 3.10 2.69 46.65 37.62 24.04 2.66 3.29 2.80
Director3D 54.32 53.33 30.94 3.25 3.43 3.05 47.79 52.81 29.31 3.36 3.67 3.20
Prometheus3D 47.46 44.32 29.15 2.84 3.12 2.66 44.73 45.85 28.57 3.20 3.36 2.98
SplatFlow 46.09 53.24 29.48 3.29 3.25 2.93 48.85 53.71 29.43 3.47 3.65 3.26
VideoRFSplat 46.52 39.50 30.13 3.12 3.24 3.09 58.19 51.71 29.76 3.58 3.63 3.30
[…]
Table 2: Quantitative results on DPG-Bench.
Method DPG-Bench
[…]
Matrix3D-omni 53.32 42.44 56.23 37.12 10.32
Director3D 66.67 64.96 60.85 45.15 22.73
Prometheus3D 45.45 48.35 55.03 33.50 9.10
SplatFlow 69.70 68.43 65.55 50.49 40.91
VideoRFSplat 36.36 56.93 66.89 48.53 31.82
[…]
Table 3: Stitching enhances NVS.
[…]
SplatFlow 19.10 0.671 0.278
VideoRFSplat 19.05 0.674 0.281
Prometheus3D 19.56 0.683 0.277
AnySplat 20.85 0.695 0.238
Hunyuan + AnySplat 21.17 0.710 0.242
SVD + AnySplat 21.48 0.720 0.218
CogVid + AnySplat 21.32 0.716 0.222
[…]
Quantitative Results. Tables 1 and 2 show the results for the three text-to-3DGS benchmarks. Notably, both tested VIST3A variants exhibit superior performance across all datasets and evaluation metrics. On T3bench, both Wan+AnySplat and Wan+MVDUSt3R consistently outperform all baselines, with particularly large margins in Imaging Quality and Coherence score. For the more complex scene-level synthesis of SceneBench, our models reach Imaging Quality scores >60 and Coherence scores >3.8, again a marked improvement over prior art. On DPG-bench, our models greatly outperform the baselines, mostly scoring >75 (often even $\approx$85), values that previously seemed out of reach. The consistent gains on T3bench, SceneBench, and DPG-bench demonstrate the effectiveness and versatility of our stitching approach for text-based 3D scene generation. We attribute these results to the power of foundational contemporary video and 3D models, which our stitching and fine-tuning scheme unlocks for the purpose of 3D generative modeling.
Qualitative Results. Figure 4 qualitatively compares VIST3A (Wan+AnySplat) to several baselines. In line with the quantitative results, VIST3A produces superior, visually compelling, and geometrically coherent renderings that closely follow the input prompts; whereas previous methods tend to exhibit artifacts, structural distortions, and poor text alignment. Further qualitative results, including Wan+MVDUst3R and Wan+AnySplat variants of VIST3A, as well as text-to-pointmap examples, can be found in Appendix E. Interestingly, we find that, even without specific training on very long image sequences, VIST3A can generate coherent large-scale scenes by extending the number of frames generated by the LDM. This demonstrates that our framework preserves the ability of video generator and the 3D decoder to handle long sequences. Examples are depicted in Fig. 13.
Figure 4: Qualitative results for 3DGS generation. We show samples from T3Bench (top), SceneBench (middle), and DPG-bench (bottom). VIST3A generates realistic and crisp 3D scenes and adheres to intricate details in the prompt.
4.3 Main Results: Model Stitching
Stitching the 3D foundation models from Section 4.1 with a video VAE yields two variants: a VAE for Gaussian splats (AnySplat + video VAE) or a VAE capable of reconstructing pointmaps and camera poses (MVDust3R or VGGT + video VAE). In the following, we evaluate both variants.
Evaluation protocol. For 3DGS models, we evaluate novel-view synthesis on RealEstate10K (Zhou et al., 2018), with 8 source and 4 target images. For 3D reconstruction models, we follow Pi3 (Wang et al., 2025d) and assess pointmap quality on ETH3D (Schöps et al., 2017), and camera pose estimation on RealEstate10K and ScanNet (Dai et al., 2017). Specifically, Accuracy (Acc.), Completion (Comp.), and Normal Consistency (N.C.) are used for pointmap estimation, while camera pose estimation is evaluated with Relative Rotation Accuracy (RRA) and Relative Translation Accuracy (RTA) at 5° and their AUC up to 30°.
Novel view synthesis. Table 3 reports results on RealEstate10K. Stitching AnySplat onto any video model always improves over using AnySplat alone. We attribute the gains to the richer appearance representation of video VAE latents. The experiment is consistent with the results of Wonderland (Liang et al., 2025), where operating in latent space rather than RGB space also benefits 3DGS. Moreover, our stitched VAEs outperform the earlier VAE-based approaches. Remarkably, we surpass Prometheus3D and VideoRFSplat despite their use of camera poses and large-scale training data, showing that stitching high-performance 3D models is indeed an effective strategy to obtain powerful 3D VAEs.
Pointmap reconstruction results. Table 4 shows that stitching preserves the accuracy and completeness of the original 3D foundation models: both pointmap quality and camera pose accuracy barely change when using video encoder latents as input. The results confirm that stitching achieves its goal, to take advantage of the pretrained models’ 3D reconstruction capabilities and repurpose them for generative modeling, without relying on large training datasets or labels.
Table 4: Results of point map reconstruction with stitched models.
Method Pointmap Estimation Camera Pose Estimation
ETH3D RealEstate10K ScanNet
[…]
Mean Med. Mean Med. Mean Med.
MVDUSt3R 0.400 0.291 0.376 0.159 0.805 0.905 98.66 12.91 42.34 0.015 0.019 0.691
VGGT 0.263 0.188 0.197 0.120 0.855 0.961 99.51 15.75 50.06 0.015 0.015 0.500
Hunyuan+MVDUSt3R 0.405 0.288 0.399 0.166 0.802 0.887 98.36 12.40 41.97 0.016 0.019 0.668
SVD+MVDUSt3R 0.410 0.310 0.387 0.168 0.804 0.899 98.12 12.67 41.69 0.016 0.020 0.690
CogVid+MVDUSt3R 0.412 0.281 0.387 0.157 0.781 0.888 98.29 12.36 41.96 0.016 0.019 0.680
[…]
4.4 Ablations
(a) log-MSE value in Eq. 2
[…]
Figure 5: MSE and pointmap quality on ETH3D vs. to stitching layer. Lower MSE in the stitching layer correlates with better 3D reconstruction.
Impact of the stitching index (Sec 3.1). We pick the best layer for stitching according to a fairly simple criterion, namely the one that best supports a linear transfer of the encoder latents. To analyze the impact of this design, we train stitched decoders for the combination (Wan+VGGT) while varying the stitching index. In Fig. 5, we see that layers with lower stitching residual indeed yield better pointmaps, supporting the MSE of the linear stitching layer as our selection criterion. Notably, early layers tend to exhibit lower MSEs. It appears that the latents are more compatible with low-level features that retain fine details.
Impact of direct reward finetuning (Sec 3.2). As shown in Appendix D.1, direct reward finetuning is more effective than a pretrained video model on its own, as well as that same model finetuned on multi-view data, with each reward component contributing to the overall performance.
Benefits of integrated vs. sequential 3D generation. In Appendix D.2, we observe that an integrated approach is more robust to noise in the latent space, which suggests it may lead to more consistent 3D reconstruction from noise in the generation process.
5 Conclusion
We have presented VIST3A, a framework for training latent diffusion models that generate 3D content from text prompts. Our key idea is to employ model stitching as a way to integrate the generative abilities of modern video models with the 3D understanding of recent feedforward 3D models. We found that this strategy indeed leads to high-quality 3D VAEs, while not requiring labeled data or massive training runs. To then align a latent-space video generator with the stitched 3D decoder it feeds into, we design a reward-based finetuning strategy. Together, these two measures yield a family of text-to-3D models with high-quality, geometrically consistent 3D outputs. In passing, they extend 3D generation to other outputs of foundational 3D models, such as pointmaps and depthmaps. More broadly, we see great potential for model stitching as a general tool to combine two or more foundational neural networks, including latent generative models, into powerful end-to-end solutions.
Appendix A Extended related works
Pipeline-based 3D generation.
A line of recent works follows a pipeline design, chaining together multiple modules and models. Typically, the first stage generates multi-view images from text or a single input image, followed by a separate reconstruction model that lifts these views into a 3D representation (Tang et al., 2024a; Xu et al., 2024b; Zhang et al., 2024b; Li et al., 2023; Park et al., 2025), with large models such as LRM (Hong et al., 2023) often used for this step. However, since the generative and reconstruction stages are trained and executed independently, errors accumulate across these parts (e.g., view inconsistency, texture flicker). Moreover, such pipeline schemes are less robust to latent-space perturbations than approaches where generation and reconstruction are performed jointly in the same latent space (see Section E).
A second category of methods (Liu et al., 2024; Yu et al., 2024b; Gao et al., 2024; Sun et al., 2024; Wang et al., 2024a) also generates multi-view images before lifting them into 3D, but replaces large pretrained reconstruction models with per-scene optimization of NeRFs (Mildenhall et al., 2021) or 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023). While this strategy avoids reliance on pretrained decoders, it remains prone to error accumulation and requires costly per-scene optimization, making inference slow and computationally expensive.
A third line of works introduces progressive expansion and refinement pipelines (Yu et al., 2024a; Ni et al., 2025; Chen et al., 2025b; Fridman et al., 2023; Feng et al., 2025; Yu et al., 2025). Some adopt iterative warping and inpainting strategies (Yu et al., 2024a; Ni et al., 2025; Fridman et al., 2023; Yu et al., 2025), while others leverage video generative models to unfold 3D scenes in a progressive manner (Chen et al., 2025b; Feng et al., 2025). Beyond these, additional works propose elaborate multi-stage pipelines that further increase complexity (Yang et al., 2025b; Ost et al., 2025). However, such designs are overly complex and suffer from slow inference.
Alignment for text-to-2D models.
Diffusion (Ham et al., 2025; Ho et al., 2020; Sohl-Dickstein et al., 2015; Song et al., 2020) and flow-matching models (Liu et al., 2023) have achieved remarkable success in 2D generation tasks across both image and video domains, and also depth estimation (Ke et al., 2025). Building upon these advances, numerous studies have explored improvements in model architectures (Lee et al., 2024; Go et al., 2023b; Park et al., 2024; 2023), loss weighting strategies (Go et al., 2023a), and timestep or noise-level sampling schemes (Kim et al., 2024). Leveraging these developments, a variety of foundational 2D generative models for images (Rombach et al., 2022) and videos (Wan et al., 2025; Yang et al., 2024b) have recently emerged.
Furthermore, recent studies have explored several strategies for aligning pretrained text-to-2D models with human preferences: (1) direct fine-tuning with scalar rewards (Clark et al., 2023; Xu et al., 2023; Prabhudesai et al., 2024; Wu et al., 2024c; Shen et al., 2025), (2) Reward Weighted Regression (RWR) (Peng et al., 2019; Lee et al., 2023), (3) Direct Preference Optimization (DPO) (Rafailov et al., 2023; Yang et al., 2024a), and (4) PPO-based policy gradients (Black et al., 2024; Fan et al., 2023; Liu et al., 2025). In this work, we adopt direct fine-tuning, which uses gradient-based feedback to align the generative model with the stitched decoder, ensuring that the resulting latents yield high-quality, 3D-consistent outputs.
Concurrent works on interchanging parts of VAEs.
Concurrently,  Chen et al. (2025a) explores replacing pretrained VAE encoders with other pretrained visual encoders that extract semantic representations. However, their approach overlooks the compatibility between the VAE latent space and the representation space of the substituted vision encoder, which consequently requires extensive retraining to achieve alignment. In contrast, we explicitly measure the similarity between the VAE latent space and the representations of each layer in 3D models, and stitch the most linearly transferable layer into the latent space. As a result, the stitched model achieves seamless integration without requiring extensive retraining.
Appendix B Methodology Details and Its Implementation
In this section, we provide additional details about the methodology behind VIST3A, extending the description given in Section 3. We first elaborate on the architectural and training aspects of our stitching method in Section B.1, including the stitching layers and loss functions used for MVDUSt3R (Tang et al., 2025), VGGT (Wang et al., 2025a), and AnySplat (Jiang et al., 2025). Section B.2 then details the direct reward finetuning methodology, outlining the reward formulations and their implementation for each 3D model (VGGT, AnySplat, and MVDUSt3R).
B.1 Model Stitching
Stitching layer.
We implement the stitching layer $\mathbf{S}$ as a single Conv3D layer. Relying only on Conv3D parameters to align the spatial and temporal dimensions between the latent and the features from $F_{k+1:l}$ can result in unnatural configurations, such as excessively large padding size. To address this, we first interpolate the latent representation to the target dimensions and then apply Conv3D, which provides a cleaner alignment of spatial and temporal dimensions. This design still admits a closed-form expression of the stitching layer, as shown in Eq. 2.
Loss function for each 3D model.
We train the stitched VAE using an $\ell_{1}$ loss between its outputs and those of the original 3D model. Since 3D model outputs often consist of multiple components, we compute the $\ell_{1}$ loss for each component separately and then aggregate them with a weighted sum. Assigning equal weights can destabilize training and even cause divergence, since some components (e.g., confidence terms) have much larger scales than others. To mitigate this, we reweight the component losses to approximately balance their scales. The specific weighting strategy is adapted to each 3D model as follows:
  • MVDUSt3R. The outputs consist of pointmaps, confidence scores for the pointmaps, and 3D Gaussian primitives We assign a weight of $10^{-2}$ to the confidence term, while pointmap and Gaussian primitive losses are left unscaled.
  • VGGT. Outputs include pointmaps, depth maps, camera poses, and confidence for both pointmaps and depth. In addition, following VGGT’s practice, we add gradient-based regularization losses on pointmaps and depth maps. We weight the pose loss by $5$, all confidence terms by $5\times 10^{-3}$, and gradient regularization losses by $5\times 10^{-3}$. Other losses remain unscaled.
  • AnySplat. Outputs include depth maps, Gaussian primitives, confidence for both depth and Gaussian primitives, camera poses, and anchor features. Additionally, we introduce gradient-based regularization losses on the depth maps. We weight all confidence terms by $10^{-2}$, gradient regularization losses by $5\times 10^{-3}$, Gaussian scale parameters by $10$, and anchor features by $0.1$. Depth and other Gaussian parameters are left unscaled.
Hyperparameters and implementation details.
For the stitching layer $\mathbf{S}$, we adopt a single 3D convolution with kernel size, stride, and padding chosen to align the latent features from the video VAE with the representation space of each 3D model:
  • MVDUSt3R: a 3D convolution with kernel size $5\times 7\times 7$, output channels $1024$, stride $1\times 3\times 3$, and padding $2\times 0\times 0$.
  • VGGT: a 3D convolution with kernel size $5\times 3\times 3$, output channels $1024$, stride $1\times 2\times 2$, and padding $2\times 1\times 1$.
  • AnySplat: a 3D convolution with kernel size $5\times 3\times 3$, output channels $1024$, stride $1\times 2\times 2$, and padding $2\times 1\times 1$.
Before applying the convolution, the interpolation layer recovers the temporal dimension compressed by the video VAE and adjusts the spatial size so that it matches the resolution expected by the feedforward 3D model. The input resolution of the video VAE is set to $384\times 384$ for MVDUSt3R and $512\times 512$ for both AnySplat and VGGT, as these configurations empirically yield stable training for the respective generative backbones. We employ LoRA with rank $r=64$ and scaling factor $a=32$ to Conv2D and linear layers across all cases.
B.2 Direct Reward Finetuning
Reward details.
We combine CLIP-based scores and HPSv2.1 human preference scores to construct rewards for both multi-view image quality and 3D representation quality. Specifically, we use DFN (Fang et al., 2024) as the CLIP model and HPSv2.1 (Wu et al., 2023). Given an image $I$ and its associated prompt $c$, we denote the HPSv2.1 score as $s_{\text{hps}}$ and the DFN CLIP score as $s_{\text{clip}}$. The quality reward is then defined as
$R_{\text{quality}}(I,c)=s_{\text{clip}}(I,c)+s_{\text{hps}}(I,c)-2,$ (4)
which implies that maximizing the reward is equivalent to maximizing the underlying score.
For the multi-view image quality reward, we compute the scores using the multi-view images decoded from the video decoder and their corresponding prompts. For the 3D representation quality reward, we compute the scores using the rendered images obtained from the 3D representation reconstructed by the stitched decoder, together with the same prompts.
[…]
Here, the negative sign ensures that maximizing the reward corresponds to minimizing both the $\ell_{1}$ distance and the perceptual discrepancy between the decoded and rendered images.
However, applying these rewards to all decoded multi-view images and their rendered counterparts is computationally expensive. To reduce computational cost, we compute all rewards only on two sampled decoded views and their corresponding rendered images. The final reward is then obtained by a weighted combination of the three components: the multi-view image quality reward and the 3D representation quality reward are each scaled by $1/16$, while the 3D consistency reward is scaled by $0.05$. These scaled terms are summed to form the overall training reward.
Alignment Algorithm. For alignment, we adopt DRTune (Wu et al., 2024c)-style direct reward finetuning, which enables stable reward optimization through selective gradient computation.
Algorithm 1 One Training Iteration of Alignment Training
1:Input: generative model $\theta$, reward $r$, sampling step range $T_{1},T_{2}]$, # of gradient enabled steps $K$, prompt $c$, data $D$.
[…]
5:Define $t$-step schedule $\{\tau_{j}\}_{j=0}^{t}$ with $\tau_{0}=T,\tau_{t}=0$
[…]
7:for $j=1$ to $t$ do $\triangleright$ Denoising from $T$ to 0
8:  $\hat{z}_{\tau_{j}}\leftarrow\texttt{stop\_grad}(z_{\tau_{j}})$
9:  if $j\in t_{\text{train}}$ then
[…]
11:  else
[…]
16:Backpropagate $\nabla_{\theta}L_{\text{total}}$, then optimize $\theta$
We outline one training iteration of our finetuning in Algorithm 1. First, we calculate the generative loss using multi-view datasets, then simulate the denoising process. Since matching the full number of inference-time denoising steps during training is costly, we instead sample $t$ steps from a reduced range $T_{1},T_{2}]$ to lower the computational burden. Additionally, to reduce time and memory costs, we only enable gradient computation at $K$ selected training steps $t_{\text{train}}$ out of the total $t$ steps. Following DRTune, the input $z_{\tau}$ to the generative model is detached at each step to stabilize optimization. Finally, we calculate the reward from the sampled latent and combine it with the generative loss by subtraction (for maximization) before backpropagation and parameter updates.
Hyperparameter in sampling process.
For generating samples required in the $T_{1},T_{2}]$ direct reward tuning stage, we set $T_{1}=10$ and $T_{2}=50$ in Algorithm 1, ensuring that the number of diffusion steps is smaller than the typical steps in inference. The number of gradient-enabled steps is set to $K=2$ to reduce memory consumption during training. For scheduling, we adopt the default scheduler from Wan 2.1 (Wan et al., 2025).
Appendix C Details on Experimental Setups
C.1 Training Setup
Setup for stitching layer search.
To identify the stitching layer, we rely on representations from the feedforward 3D model and the corresponding latents computed on the same dataset. Specifically, we utilize a subset of the DL3DV dataset, comprising 200 scenes for VGGT, 800 scenes for AnySplat, and 3,200 scenes for MVDUSt3R, with only 13 views per scene used for the search. We limit our search to the encoder layers of each model, as we observe that MSE values consistently increase within deeper layer indices.
Setup for stitched VAE finetuning.
We train on a combination of the DL3DV and ScanNet datasets, defining one epoch as a full pass over DL3DV and two passes over ScanNet. For each training iteration, a number of scenes are sampled according to the batch size. From each selected scene, we randomly sample 9 or 13 views to serve as input samples for training. The models are trained for 50 epochs in total. The batch sizes are set to 12 for VGGT, 24 for MVDUSt3R, and 12 for AnySplat. The learning rate is fixed at $2\times 10^{-4}$ for all models with cosine decay scheduling and 500-step warmup. For training, we use AdamW (Loshchilov & Hutter, 2017), apply gradient clipping with a norm threshold of 1.0, and use gradient checkpointing on each stitched VAE block to reduce memory consumption. In addition to LoRA parameters, for AnySplat and VGGT, we also finetune register tokens and class tokens. This is necessary because we remove the earlier layers that originally process these tokens into intermediate representations, requiring adaptation of the token handling mechanism. We further utilize gradient checkpointing for every stitched VAE block.
Setup for generative model finetuning.
We finetune the generative models using only the DL3DV dataset. For generative loss computation, we use a batch size of 12 with 13 views per scene. Reward calculation uses a prompt batch size of 4, with 13 views for AnySplat and MVDUSt3R, and 9 views for VGGT. We again adopt AdamW with a learning rate of $1\times 10^{-4}$, apply gradient clipping at a 0.1 norm, and train LoRA parameters with rank 8 and alpha 16. Gradient checkpointing is enabled for all model blocks to reduce memory usage.
C.2 Detailed Evaluation Protocol
Details for 3D generation evaluation.
For T3Bench, we evaluate on all 300 prompts, in contrast to prior works that considered only the 100 single-object-with-surroundings subset. SceneBench is evaluated on 80 prompts from the Prometheus3D (Yang et al., 2025c) prompt set, targeting scene-level generation. For DPG-Bench, we sample 100 prompts from the original 1K-prompt dataset.
For Matrix3D-omni, we used their official code for text-to-generation and employed Panorama LRM for reconstruction during inference. For SDS-based methods like SplatFlow and Director3D that perform refinement, we evaluated the final results after SDS optimization. We generate 13 frames for all models using 80 denoising steps, and apply classifier-free guidance (Ho & Salimans, 2022) with a scale of 7.5. We observed that the Gaussian splatting produced by the MVDUSt3R model does not generalize well across diverse domains, often failing to estimate the scale of primitives. To address this issue, we refined the Gaussian primitives using the source view for 100 optimization steps, minimizing a reconstruction loss defined as MSE + 0.05 × LPIPS. For this refinement, we used the Adam optimizer with separate learning rates for each parameter group: 2e-4 for means, 5e-4 for opacity, 5e-4 for scale, 1e-4 for rotation, and 0 for rgbs. This lightweight refinement effectively corrected the scale estimation errors. For our text-to-3DGS evaluation, we render 8 random viewpoints from the generated Gaussian Splatting representations for assessment.
We evaluate our method and baselines across a range of metrics. To measure the semantic similarity between the input prompt and the rendered images of the generated 3DGS, we compute the CLIP score using the clip-vit-base-patch16 model. Additionally, we adopt the VBench (Huang et al., 2024) framework to assess key image properties. For Imaging Quality, which targets low-level distortions, we employ the same MUSIQ model (Ke et al., 2021) in VBench. For Aesthetic Quality, we use the LAION aesthetic predictor to evaluate the color richness and artistic merits, again following VBench. The predictor’s native 0-10 rating is linearly normalized to a 0-1 scale for our analysis.
For a more comprehensive assessment of generative quality, we utilize the Unified Reward model (Wang et al., 2025c), which is based on the powerful Qwen 2.5-7B Vision Language Model (Team, 2025) [^2] https://huggingface.co/CodeGoat24/UnifiedReward-qwen-7b. This provides fine-grained, pointwise scores on complex attributes equipped with a powerful understanding capability. By feeding the input prompt and rendered images into a format adapted from the official implementation script [^3] https://github.com/CodeGoat24/UnifiedReward/blob/main/inference_qwen/image_generation/qwen_point_score_ACS_image_generation.py, we obtain scores for three key aspects:
  • Alignment: How well the image content matches the text prompt.
  • Coherence: The logical and visual consistency of the image, free of distortions.
  • Style: The aesthetic appeal of the image, independent of prompt accuracy.
This suite of metrics enables a robust and multifaceted evaluation of our model’s performance.
Details for model stitching evaluation.
For novel-view synthesis, we follow prior works (Go et al., 2025a; b) and adopt an 8-frame input setup to evaluate performance on 4 target views. To accommodate the fixed-length input requirements of video VAE architectures due to temporal compression, we pad shorter sequences by duplicating the final frame. For estimating the camera poses of the target views, we adopt the strategy from AnySplat (Jiang et al., 2025), which jointly predicts the poses and renders the corresponding images. This contrasts with previous VAE-based methods that presume access to ground-truth camera poses for rendering.
For pointmap and camera pose estimation evaluation, we use a 13-frame input setup. Since our stitched VAE’s encoder is a video VAE, we arrange the multi-view images (typically provided unordered by previous works) into sequences with smooth view transitions to resemble video input. We adopt Pi3 (Wang et al., 2025d) official evaluation code and follow their preprocessing pipeline.
Appendix D Further Ablation Studies
D.1 Impact of Direct Reward Finetuning
In the following, we conduct an ablation study to analyze the effects of our direct reward finetuning, comparing our full method against four well-defined baselines:
  • (1) Finetuning-free: Here, we use the original pretrained video model. Since our finetuning freezes the encoder, its latent space remains compatible with our 3D stitched decoder.
  • (2) Multi-view Only: The model finetuned with only the flow-matching loss on multi-view data, serving as our primary baseline before rewards are introduced.
  • (3) Multi-view + Consistency: The model finetuned with both the multi-view loss and the 3D-consistency reward. This isolates the impact on the 3D consistency reward.
  • (4) Multi-view + Quality: The model finetuned with both the multi-view loss and the quality reward. This isolates its impact on quality reward.
To ensure a fair comparison against reward-based methods, which often take more time for one training iteration, the finetuning variant on multi-view data was trained for the same wall-clock duration.
Table 5: Ablation study on direct reward finetuning on SceneBench. We compare (1) no finetuning; (2) multi-view-only finetuning (generative loss only); (3) reward tuning with 3D-consistency reward only; (4) reward tuning with quality reward only; and (5) reward tuning with both rewards (full).
Method Imaging Aesthetic CLIP Unified Reward
Alignment Coherent Style
Finetuning-free 50.56 53.70 28.14 3.101 3.354 3.393
Multi-view only 54.56 52.08 29.71 3.622 3.834 3.351
Multi-view + 3D Consistency 38.67 50.59 29.77 3.581 3.767 3.275
Multi-view + Quality 62.27 58.23 30.34 3.643 3.842 3.358
[…]
Table 5 reports the quantitative results. The finetuning-free baseline yields the lowest performance. Lacking any 3D-aware training, it frequently produces geometrically inconsistent outputs and suffers from significant visual artifacts when its native resolution is adapted to our 3D decoder. Introducing multi-view supervision (Multi-view Only) substantially improves 3D consistency and overall performance, confirming the value of this training signal.
The reward components have distinct effects when added to the multi-view objective. Training with the 3D-consistency reward (Multi-view + Consistency) leads to a notable performance drop, as the model optimizes for geometric correctness at the expense of detail, resulting in overly blurred images. Conversely, adding the quality reward (Multi-view + Quality) achieves substantial improvements across most metrics by enhancing prompt coherence and aesthetic appeal.
Finally, our full method, which combines both rewards with multi-view training, achieves the best imaging quality and Unified Reward scores. While its aesthetic and CLIP scores are slightly below the Multi-view + Quality variant, the marked improvement in imaging quality demonstrates that our combined objective successfully guides the model to generate visually sharp and geometrically faithful 3D representations.
D.2 Benefits of integrated vs. sequential 3D generation
In our model-stitching design, generation and reconstruction take place in the shared latent space of the video diffusion VAE and the stitched 3D decoder. A common alternative is a sequential pipeline that decodes latents into RGB frames before applying a feedforward 3D model (e.g., VGGT) without further adaptation. To probe the core benefit of our unified formulation, we injected controlled perturbations into the latent representation, using
$z^{\prime}=z+\alpha\,\|z\|\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),$ (6)
where $\alpha$ is a scalar controlling the perturbation strength. We then compared two paths: (i) decode the corrupted latent to RGB and feed the images sequentially into the original VGGT (baseline), and (ii) directly input the noised latent into our stitched 3D decoder (unified latent framework).
(a) Acc. $\downarrow$
[…]
(d) Reconstructed images through VAE according to $\alpha$
Figure 6: Pointmap estimation performance comparison on ETH3D dataset between the stitched VGGT and the sequential approach (VAE followed by VGGT) under varying noise scales injected into the latent space. The stitched model demonstrates greater robustness to noise injection in the VAE.
Figure 6 reports pointmap estimation performance on ETH3D as a function of noise level $\alpha$. Our stitched VGGT consistently outperforms the sequential decode-and-reconstruct pipeline under noise injection, indicating that the VAE decoder in the sequential path amplifies errors. Moreover, as shown in Fig. 6(d), the performance gap is observed even at noise levels ($\alpha=1e^{-4}$ to $2e^{-4}$) where visual artifacts are hardly perceptible. This suggests that the unified design offers stronger robustness, as imperceptible perturbations from the noise of generative processes can already degrade the sequential pipeline.
Appendix E Additional Experimental Results
To comprehensively validate each component of VIST3A, we present additional experiments in this section.
(a) Wan
(b) Hunyuan
(c) SVD
(d) CogvideoX
Figure 7: Log-MSE values in Eq. 2 across various video VAEs. Early layers of feedforward models show lower MSE values within each VAE architecture. While lower MSE correlates with better stitching performance within the same VAE (e.g., layer 2 outperforms layer 16 for Wan in Fig 5), absolute MSE values cannot predict performance across different VAE architectures. For instance, despite CogVideoX and Hunyuan + AnySplat having the lowest absolute MSE (0.008), SVD + AnySplat achieves the best performance (21.48 PSNR) in Table 3.
Analysis on searched stitching index.
In Section 4.4, we showed that earlier layers in the network tend to be more linearly correspondent. We extend this analysis to various VAE architectures, including CogVideoX, SVD, Hunyuan, and Wan, paired with MVDUSt3R and AnySplat, to observe the generalizability of this finding.
Figure 7 shows the log-MSE values measuring linear transferability between latents and the feedforward 3D model’s representations. From the results, early layers of 3d models consistently show lower MSE values across all VAE-feedforward 3D model combinations. This supports the hypothesis that latent representations capture low-level features for input reconstruction, which are more linearly transferable to the early layers of the feedforward 3D model that also encode such features. However, the results reveal an important distinction: while relative MSE ordering within each VAE architecture correlates with stitching performance (as in Section 4.4), absolute MSE values across different VAEs do not predict cross-architecture performance. For instance, CogVideoX + AnySplat achieves the lowest absolute MSE (0.008) but delivers 21.32 PSNR in Table 3, while SVD + AnySplat with a higher MSE (0.012) achieves superior performance at 21.48 PSNR. This indicates that optimal stitching layers must be identified independently for each VAE-3D model pair.
Additional qualitative results.
We present additional qualitative results of VIST3A with Wan + AnySplat in Fig. 810. Text-to-pointmap generation results obtained by combining VGGT with Wan through VIST3A are shown in Fig. 11. Finally, Fig. 12 illustrates VIST3A results with MVDust3R + Wan.
Figure 8: Qualitative comparison of 3DGS generation. The top two rows show samples from DPG-Bench, and the bottom two rows present samples from T3Bench. VIST3A generates realistic scenes with fine-grained details that faithfully reflect the input prompt, outperforming baselines.
Figure 9: Qualitative comparison of 3DGS generation on SceneBench. VIST3A outperforms baselines by generating higher-fidelity scenes with accurate geometry and appearance.
Figure 10: Generated 3D scenes from VIST3A: Wan + AnySplat. These are 3DGS viewed directly in the interactive viewer. VIST3A preserves high visual quality even under noticeably altered camera trajectories, demonstrating robustness and stability across novel viewpoints.
Figure 11: Qualitative results on text-to-poinmap generation. By integrating VGGT, VIST3A generates structurally consistent pointmaps and fine-grained details across diverse prompts.
Figure 12: Qualitative comparison of 3DGS generation on SceneBench - VISTA: Wan+MVDUSt3R.
Figure 13: Generated 3D scenes from VIST3A: Wan + AnySplat bt extending the number of frames. These are 3DGS viewed directly in the interactive viewer. VIST3A preserves high visual quality even under noticeably altered camera trajectories, demonstrating robustness and stability across novel viewpoints.
Appendix F Limitations
While our approach demonstrates strong results, it also has certain limitations. Our stitched model inherits its encoder from a video generation model, which is inherently designed for sequential, temporally coherent video input. Consequently, its performance is not guaranteed for arbitrarily unordered inputs, such as typical multi-view image datasets. To ensure the encoder operates effectively, the input images must be arranged into a coherent sequence that simulates the smooth view transitions of a video clip.
Appendix G Use of Large Language Models
LLMs were used exclusively for text polishing and grammar refinement.